OpenStudy (anonymous):

Transformations??? I don't understand State the various transformations applied to the base function f(x)=|x| to obtain a graph of the functiong(x) = |x + 1| - 2.

OpenStudy (anonymous):

(a.)Horizontal shift of 1 unit to the left and a vertical shift upward of 2 units. (b.)Horizontal shift of 1 unit to the left and a vertical shift downward of 2 units. (c.)Horizontal shift of 1 unit to the right and a vertical shift downward of 2 units. (d.)Horizontal shift of 1 unit to the right and a vertical shift upward of 2 units.

OpenStudy (solomonzelman):

SHIFTS: \(\large\color{black}{ f(x)=\left| x \right| ~~~~~\bf{\rightarrow}~~~~~f(x)=\left| x \color{blue}{ -~\rm{c} }\right| }\) \(\large\color{blue}{ ~\rm{c} }\) units to the right. \(\large\color{black}{ f(x)=\left| x \right| ~~~~~\bf{\rightarrow}~~~~~f(x)=\left| x \color{blue}{ +~\rm{c} }\right| }\) \(\large\color{blue}{ ~\rm{c} }\) units to the left. \(\large\color{black}{ f(x)=\left| x \right| ~~~~~\bf{\rightarrow}~~~~~f(x)=\left| x \right| \color{blue}{ +~\rm{c} }}\) \(\large\color{blue}{ ~\rm{c} }\) units to up. \(\large\color{black}{ f(x)=\left| x \right| ~~~~~\bf{\rightarrow}~~~~~f(x)=\left| x \right| \color{blue}{ -~\rm{c} }}\) \(\large\color{blue}{ ~\rm{c} }\) units to down.

OpenStudy (solomonzelman):

your function is supposed to be, \(\large\color{black}{ f(x)=\left| x +1\right|+2 }\) is that correct?

OpenStudy (solomonzelman):

or g(x)... f(x), g(x), h(x) ... makes no difference here.

OpenStudy (anonymous):

it's supposed to be f(x) = |x + 1| - 2, not add 2 I messed up something else though, forgot to include the base function in original question, changing that now

OpenStudy (solomonzelman):

\(\large\color{black}{ f(x)=\left| x +1\right|-2 }\)

OpenStudy (solomonzelman):

sure...

OpenStudy (anonymous):

just added the base function to original question, and changed the equals to subtract 2

OpenStudy (solomonzelman):

sure:)

OpenStudy (solomonzelman):

Do you understand tyhe rules I posted?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wait hang on they're starting to make sense

OpenStudy (anonymous):

so the units will move to the left?

OpenStudy (solomonzelman):

how many units to the left? which one is the shift to the left, the +1 on the inside, or the -2 on the outside?

OpenStudy (anonymous):

1 unit to the left, the shift to the left is the +1 on the outside

OpenStudy (anonymous):

so far my answer choices are either a or b?

OpenStudy (solomonzelman):

yes that is correct. the +1, moves the f(x) 1 unit left.

OpenStudy (solomonzelman):

Now, the "-2" part on the outside, is a shift where, and by how many units?

OpenStudy (anonymous):

it's a downward shift because subtracting downward on a coordinate grid? and 2 units

OpenStudy (solomonzelman):

yes. you are going 2 units down.

OpenStudy (anonymous):

So my answer is B, Horizontal shift of 1 unit to the left and a vertical shift downward of 2 units.?

OpenStudy (solomonzelman):

So g(x), the new function, is 1 units to the left and 2 units down, from the (initial) f(x). yes, B....

OpenStudy (anonymous):

thank you!!

OpenStudy (solomonzelman):

I would write these rules somewhere on a paper so that you rememeber them. Because you will need them ....

OpenStudy (anonymous):

i will

OpenStudy (solomonzelman):

not only in algebra, but even on in calc1:) .... you welcome!